2y-x=-8 3x+4y=-6
I have to solve this by elimination and Im really confused. Could someone help please?

Answer by checkley71(8403)

2Y-X=-8
4Y+3X=-6
MULTIPLY THE FIRST EQUATION BY 3 AND SUBTRACT THEM.
6Y-3X=-24
4Y+3X=-6
---------------------------------
10Y=-30
Y=-30/10
Y=-3 ANSWER.
2*-3-X=-8
-6-X=-8
-X=-8+6
-X=-2
X=2 ANSWER.
PROOF
4*-3+3*2=-6
-12+6=-6
-6=-6